OpenStudy (theviper):

\(\Large{TUTORIAL:-}\) How to find the quadratic & cubic polynomials if there zeroes are given .....

OpenStudy (anonymous):

(x-a)(x-b)....... where a,b and so on are the roots

OpenStudy (vishweshshrimali5):

@Brittney.Anne ask ur questions separately

OpenStudy (theviper):

QUADRATIC polynomial :- If zeroes are \(\alpha\ \&\ \beta \) then polynomial will be find by this :- \[x^2-x(\alpha + \beta)+\alpha \beta\] CUBIC polynomial :- If zeroes are \(\alpha, \beta \ \& \ \gamma \) then polynomial will be find by this :- \[x^3-x^2(\alpha+\beta+\gamma)+x(\alpha \beta+\beta \gamma+\gamma \alpha)-\alpha \beta \gamma\]

OpenStudy (vishweshshrimali5):

Ye tutorial tha ............ viper ?

OpenStudy (vishweshshrimali5):

:D

OpenStudy (vishweshshrimali5):

ADVICE : never post such tutorials again ;)

OpenStudy (vishweshshrimali5):

Just joking

OpenStudy (vishweshshrimali5):

:D:

OpenStudy (vishweshshrimali5):

;)

mathslover (mathslover):

too much comments(off topic) 1) friends delete your off topic comments that may decrease the publicity (TRP) of @TheViper 2) short and sweet tutorial... good @TheViper but please explain more or give examples

OpenStudy (unklerhaukus):

neat tutorial

mathslover (mathslover):

:D @UnkleRhaukus

OpenStudy (jiteshmeghwal9):

\[\LARGE{\color{red}{\cal{too \space short \space tutorial \space bro ;)}}}\]

OpenStudy (jiteshmeghwal9):

just joking :)

OpenStudy (unklerhaukus):

\[:\ddot D: \]

OpenStudy (lgbasallote):

that is what's called evolutiion @mathslover ...tutorials are short and less detailed but more notations nowadays >:))

mathslover (mathslover):

no comments...

OpenStudy (vishweshshrimali5):

@TheViper see the irony............. Your lattice of question say \Large{TUTORIAL:-}\ ............. and your tutorial is small(but gud)............ Just noticed it today

OpenStudy (jiteshmeghwal9):

LOL ;)

OpenStudy (vishweshshrimali5):

;)